OpenStudy (anonymous):

Vijay has an annual salary of 47,000, and his company pays him twice a month. What is the gross income in each paycheck that Vijay receives?

OpenStudy (anonymous):

The company pays every 2 weeks and we need to find what the company pays every 2 weeks. \[\frac{ 54weeks }{ $47000 }=\frac{ 2weeks }{ x }\] A year has 54 weeks and we\re going to use weeks because we need to know the value of each paycheck. Now cross multiply it \[54weeks * (x)=47000*2weeks\] now solve for x \[x=94000/54weeks =1740.74\] therefore the value of each paycheck is $1740.74

OpenStudy (anonymous):

let me know if you didn't understand anything i'll be back

OpenStudy (anonymous):

if you understand it please give me a medal